OpenStudy (anonymous):

A disgruntled fast food operator hurls and Egg Mcmuffin at 18m/s, 26degrees above the horizontal offa 12 meter high cliff. what are its horizontal and vertical velocities? How long will it be in the air? How far from the cliff's base will it land?

OpenStudy (dumbcow):

similar to last one...the horizontal and vertical velocities are 18*cos(26) 18*sin(26) \[h(t) = -4.9t^{2}+18\sin(26)t +12\] \[x(t) = 18 \cos(26) t\]

OpenStudy (anonymous):

oohh thanks man was confused a lil a bit off in my thinking today...

OpenStudy (dumbcow):

no prob :)